# Federated Instruction Tuning of LLMs with Domain Coverage Augmentation

### v3

@@ 1 Introduction
Recently, federated instruction tuning (FedIT) has gained attention as a novel approach that leverages the principles of federated learning (FL) to facilitate collaborative training of large language models (LLM) in distributed environments while maintaining the confidentiality of private data (McMahan et al., 2017; Ye et al., 2024b). This methodology allows for the exchange of model parameters among distributed data holders, thereby achieving a careful balance between privacy preservation and efficient model optimization. Despite the establishment of various FedIT frameworks (Ye et al., 2024b; Kuang et al., 2023; Zhang et al., 2024a), existing literature has not adequately addressed the practical challenges that Federated Domain-specific Instruction Tuning (FedDIT) may encounter in real-world applications. For instance, FedIT generally necessitates a sufficient amount of instruction data for fine-tuning, which is often a shortage in domain-specific fine-tuning contexts (Zhang et al., 2024d).
@@
Additionally, the factors affecting FedDIT are still unclear. Compounding this uncertainty, introducing augmented instructions may further complicate results, making it difficult to ascertain effective improvement strategies. Shepherd (Zhang et al., 2024a) approaches this problem from the perspective of heterogeneity, constructing heterogeneity based on the topic of general instruction datasets. It demonstrates that, unlike the consensus of traditional FL, for general FedIT, heterogeneity has a positive effect. By aggregating diverse instructions from clients, the diversity increases, thereby enhancing the model’s adaptability to various tasks. However, it just scratches the surface and does not explore issues in FedDIT.
@@
We demonstrate the effectiveness of FedDCA through comprehensive experiments conducted across four domains: code, medical, financial, and mathematical. These are compared against a range of baselines, which can be categorized into unaugmented and augmented methods. In the unaugmented setting, our method is compared with FedIT, which includes four orthodox FL techniques: FedAvg (McMahan et al., 2017), FedProx (Li et al., 2020), SCAFFOLD (pmlr-v119-karimireddy20a), and FedAvgM (Hsu et al., 2019). In the augmented setting, we compare FedDCA against methods such as random sampling, direct retrieval, LESS (Xia et al., 2024), and Self-Instruction (Wang et al., 2023c). Additionally, we present the performance outcomes of FedDCA when applied under various FL strategies. We also compare the computational efficiency on the client side between FedDCA and its variant, FedDCA. For privacy analysis, experiments against memory extraction attacks are conducted to evaluate how different quantities of retrieved public data affect the privacy of client local data. Results indicate that while reliance on local data increases memorization of sensitive information, the risk of privacy leakage diminishes or converges in the augmented setting as the training rounds progress.
@@ 2 Preliminaries
FedDIT aims to leverage cross-client private domain-specific instruction data and utilize the multi-domain public data on the server to achieve instruction augmentation, collaboratively enhancing the model’s performance in specific domains. Consider $N$ distributed clients, each with local private data $D_{k}^{l}$. The server maintains a public dataset $D^{p}$ that encompasses multiple domains and is responsible for implementing data augmentation strategies and aggregating model parameters received from clients. The training process follows the standard FL protocol (McMahan et al., 2017). For computational efficiency, we adopt Low-Rank Adaption (LoRA) (Hu et al., 2021) as the fine-tuning method, which involves tuning additional parameters $\Delta\phi$ while keeping the pre-trained LLM’s parameters $\phi$ frozen. In the initial training phase, the server dispatches $\phi$ and $\Delta\phi$ to each client for $\mathcal{R}$ training rounds. In the $t$-th round, the server sends the aggregated $\Delta\phi^{t}$ to clients. Clients use $\Delta\phi^{t}$ to update their local LoRA parameters $\Delta\phi_{k}^{t}$ and conduct instruction tuning based on augmented instruction datasets $D_{k}$. Subsequently, the clients return $\Delta\phi_{k}^{t+1}$ to the server. The server then aggregates $\{\Delta\phi_{k}^{t+1}\mid k=1,\dots,N\}$ to obtain $\Delta\phi^{t+1}$ for the next round.
@@ 3 Problem Formulation
For better understanding, we list the frequently used notation in Appendix A.11. The objective of FedDIT is to enhance the domain-specific performance of LLMs through FL without sharing private data (Ye et al., 2024b; Zhang et al., 2024d). Under the FL framework, suppose we have $N$ clients, where each client $c_{k}$ has a local dataset $D_{k}^{l}$ with its size $N_{k}^{l}$, and an augmented dataset $D_{k}^{g}$ from the server public dataset $D^{p}$, respectively. Due to constraints such as memory, computational overhead, and maximum tolerated training time, client $c_{k}$ can accept at most $N_{k}^{p}$ public instructions. Denote the augmentation strategy as $\Lambda$, through which the server performs instruction augmentation on the public dataset. If $\Lambda$ is null, it indicates that clients conduct FedDIT solely based on their local private data, which may lead to performance degradation. Conversely, $\Lambda$ may be classified into two categories: (1) focusing exclusively on the client’s own local data distribution or (2) considering the cross-client data distribution. In conclusion, the global objective of FedDIT is defined as follows:
@@ 4.1 Optimization Problem
Suppose there exists a metric space $\mathcal{X}$, a set of cross-client local instruction embeddings $\mathcal{E}\in\mathcal{X}$, and a set of cluster centers $\mathcal{P}\in\mathcal{X}$. Each client $c_{k}$ has maximum $\xi$ clusters obtained by k-means algorithm (Vardakas & Likas, 2023). In the federated setting, communication cost is always a critical factor. Consequently, we formulate the optimization problem to include communication costs as follows:
@@ 4.4 Discussions
Computation. Through greedy client center selection, FedDCA solves the optimization problem defined in Eq. 3 in the polynomial time and is quite efficient. On the client side, the k-means algorithm (Vardakas & Likas, 2023) is $\mathcal{O}(|D_{k}^{l}|)$. On the server side, as mentioned in Section 4.1, the time complexity of greedy client center selection is $\mathcal{O}\left(N\left(N\xi+\xi\log\xi\right)\right)$. Specifically, for each client, the similarity computation between selected client center set $\mathcal{P}$ and client $c_{k}$’s cluster $\mathcal{C}_{k}$ is $\mathcal{O}\left(N\xi\right)$, and the sorting process is $\mathcal{O}\left(\xi\log\xi\right)$. For domain data retrieval, given that the sorting algorithm is $\mathcal{O}\left(\mathcal{N}\log\mathcal{N}\right)$, where $\mathcal{N}$ is the size of the public dataset. Thus the time complexity is $\mathcal{O}\left(N(\mathcal{N}\log\mathcal{N})\right)$. In addition, the heterogeneous encoder setting FedDCA further reduces the computation overhead on the client side.
@@
Privacy. The proposed method, FedDCA, does not violate the client’s privacy. Compared with other FedIT methods (Zhang et al., 2024d; Ye et al., 2024b), the difference lies in the greedy client center selection. In this stage, the client only uploads the cluster center to the server, which is the average of embeddings to its cluster. In addition, the potential privacy leakage can be further avoided through homomorphic encryption (Acar et al., 2018), which allows the server to directly compute on ciphertext for matrix multiplication for dense retrieval.
@@ 5.1 Experimental Setup
Baselines. We compare FedDCA and FedDCA with the following baselines: 1) Unaugmented methods, including zero-shot inference and FedIT, which are composed of four widely used FL methods, including FedAvg (McMahan et al., 2017), FedProx (Li et al., 2020), SCAFFOLD (pmlr-v119-karimireddy20a) and FedAvgM (Hsu et al., 2019). 2) Augmented methods, which include random sampling, direct retrieval, LESS (Xia et al., 2024), and Self-Instruct (Wang et al., 2023c). Additionally, we report FedDCA’s performance with different FL strategies in Table 2. Specifically, zero-shot inference shows the performance of the pre-trained LLM without FedDIT, which gives the lower performance bound. Direct retrieval is described in Appendix A.3. Self-Instruct augments instruction in a generative way through GPT-3.5-turbo (Sun et al., 2023). Based on the prompt provided by Self-Instruct, we define the prompts for generating instructions and responses as shown in Appendix A.7.
@@ 5.2 Performance Analysis
Efficiency. We see that direct retrieval is the best baseline in the average performance. However, it does not consider cross-client domain coverage, resulting in an overlapping retrieved data distribution. Self-Instruct (Wang et al., 2023c) represents the upper limit performance of the generation-based method, which is restricted to the high expense of cost, limited API call frequency, and heavy quality screening process. On the other hand, FewFedPIT (Zhang et al., 2024d) attempts to augment local data by leveraging the pre-trained Llama2 model to perform self-instructed generation during training. However, two concerns exist: 1) The pre-trained LLM cannot provide effective, stable, and high-quality instruction generation. 2) Generating instructions locally is very costly regarding both time and computational resources. Overall, generating instructions self-instructively is not a satisfactory method for the FedDIT scenario. LESS (Xia et al., 2024) represents the augmentation methods through gradient feature retrieval. However, its performance is underwhelming. Compounding this issue, LESS presupposes access to the validation set, which is not always available, and necessitates an initial warmup on the public dataset before calculating gradients for both the public and validation data. This process becomes computationally burdensome with large public datasets. In conclusion, FedDCA achieves better performance while requiring lower computational resources and a more relaxed training condition, which shows its efficiency.
@@ 5.4 Privacy Analysis
Next, we evaluate the privacy-preserving capability of different ratios of public data against memory extraction attacks (Carlini et al., 2021; Zhang et al., 2024c), which utilizes the autoregression nature of LLM to extract information from the memory (Xu et al., 2024).
@@
Retrieval Num Code Med. Fin. Math.
@@ A.1 Related Work
Federated Instruction Tuning. Instruction tuning has been widely applied across various application areas of large language models (LLM), serving as a key technique to enhance the capabilities and controllability of LLM (Zhang et al., 2024b; Wei et al., 2022). Recently, federated instruction tuning (FedIT) has emerged as an effective strategy for the distributed optimization of LLMs, leveraging federated learning (FL) protocols to improve the handling of privacy-sensitive tasks in real-world scenarios. So far, several FedIT frameworks (Ye et al., 2024b; a) have been established to evaluate the effectiveness of FedIT across multiple datasets, tasks, and FL methods. While these platforms provide a foundation for research, they have not yet introduced more complex federated algorithms and deeply investigate the challenging problems and factors affecting FedIT, which are crucial for advancing this field.
@@
PrivateLoRA (Wang et al., 2023b) addresses privacy and efficiency issues by exploiting the low-rank properties of residual activations to reduce communication costs, significantly lowering the communication overhead through collaborative computation between the server and clients while effectively maintaining the privacy of local data. While FewFedPIT (Zhang et al., 2024d) focuses on the few-shot learning setting in FedIT, using self-generated data by pre-trained LLMs locally to mitigate the paucity of data and first discussing memory extraction attacks within FedIT.
@@
Domain Instruction Augmentation. In the real world, there is an urgent need for training LLMs with specific functionalities (e.g., reasoning capabilities) or domain-specific LLMs (e.g., code (Nijkamp et al., 2023; Luo et al., 2023b), medical (2024-Zhang-AlpaCare-), financial (Yang et al., 2023b; a; Zhang et al., 2023a; Wu et al., 2023), mathematical (Yue et al., 2024; Luo et al., 2023a)). Existing works tend to use open-source domain-specific instruction tuning datasets for training. However, the target domain may not always have corresponding ready-made domain-specific instruction datasets. Even if they exist, these datasets are often limited in scale.
@@
Several studies investigate domain-specific instruction augmentation, which can be categorized into three aspects: 1) Reusing human-curated public datasets (Wang et al., 2023a; Zhang et al., 2023c; Jiao et al., 2023; Lee et al., 2024; Xia et al., 2024). For instance, Parrot (Jiao et al., 2023) enhances translation capabilities of LLMs in chat by converting bilingual sentences into instruction-following datasets. Furthermore, works like INSTA (Lee et al., 2024) and LESS (Xia et al., 2024) attempt efficient domain-specific instruction augmentation via dense retrieval. INSTA (Lee et al., 2024) uses instructions without responses for effective retrieval. LESS (Xia et al., 2024) assumes access to a validation set and uses the warmup LLM’s gradients of the train and validation sets for retrieval. 2) Using seed tasks for self-instruct (Luo et al., 2023b; Wan et al., 2023): Explore-Instruct (Wan et al., 2023), for example, employs activate exploration to tree-model domain tasks from both depth and breadth, increasing seed instructions’ coverage of the domain, and subsequently generating broader in-domain instructions through self-instruct. 3) Scaling instructions from the web: Recent works (Yue et al., 2024; Zhou et al., 2024) highlight the immense potential of mining naturally occurring instructions from the internet. Compared to generated data, web-mined instructions exhibit less bias and greater diversity. MAmmoTH2 (Yue et al., 2024) firstly retrieves domain-relevant texts and employs a LLM to extract Q-A pairs, further refining them into instruction-response pairs. Jiuzhang3.0 (Zhou et al., 2024) distills GPT’s instruction generation capabilities into a smaller model and then uses it to generate instructions from the internet. In conclusion, models fine-tuned with augmented instructions have shown promising domain-specific capabilities.

### v4

@@ 1 Introduction
Recently, federated instruction tuning (FedIT) has gained attention as a novel approach that leverages the principles of federated learning (FL) to facilitate collaborative training of large language models (LLM) in distributed environments while maintaining the confidentiality of private data (McMahan et al., 2017; Ye et al., 2024b). This methodology allows for the exchange of model parameters among distributed data holders, thereby achieving a careful balance between privacy preservation and efficient model optimization. Despite the establishment of various FedIT frameworks (Ye et al., 2024b; Kuang et al., 2023; Zhang et al., 2023c), existing literature has not adequately addressed the practical challenges that Federated Domain-specific Instruction Tuning (FedDIT) may encounter in real-world applications. For instance, FedIT generally necessitates a sufficient amount of instruction data for fine-tuning, which is often a shortage in domain-specific fine-tuning contexts (Zhang et al., 2024b).
@@
Additionally, the factors affecting FedDIT are still unclear. Compounding this uncertainty, introducing augmented instructions may further complicate results, making it difficult to ascertain effective improvement strategies. Shepherd (Zhang et al., 2023c) approaches this problem from the perspective of heterogeneity, constructing heterogeneity based on the topic of general instruction datasets. It demonstrates that, unlike the consensus of traditional FL, for general FedIT, heterogeneity has a positive effect. By aggregating diverse instructions from clients, the diversity increases, thereby enhancing the model’s adaptability to various tasks. However, it just scratches the surface and does not explore issues in FedDIT.
@@
We demonstrate the effectiveness of FedDCA through comprehensive experiments conducted across four domains: code, medical, financial, and mathematical. These are compared against a range of baselines, which can be categorized into unaugmented and augmented methods. In the unaugmented setting, our method is compared with FedIT, which includes four orthodox FL techniques: FedAvg (McMahan et al., 2017), FedProx (Li et al., 2020), SCAFFOLD (Karimireddy et al., 2020), and FedAvgM (Hsu et al., 2019). In the augmented setting, we compare FedDCA against methods such as random sampling, direct retrieval, LESS (Xia et al., 2024), and Self-Instruction (Wang et al., 2022). Additionally, we present the performance outcomes of FedDCA when applied under various FL strategies. We also compare the computational efficiency on the client side between FedDCA and its variant, FedDCA. For privacy analysis, experiments against memory extraction attacks are conducted to evaluate how different quantities of retrieved public data affect the privacy of client local data. Results indicate that while reliance on local data increases memorization of sensitive information, the risk of privacy leakage diminishes or converges in the augmented setting as the training rounds progress.
@@ 2 Preliminaries
FedDIT aims to leverage cross-client private domain-specific instruction data and utilize the multi-domain public data on the server to achieve instruction augmentation, collaboratively enhancing the model’s performance in specific domains. Consider $N$ distributed clients, each with local private data $D_{k}^{l}$. The server maintains a public dataset $D^{p}$ that encompasses multiple domains and is responsible for implementing data augmentation strategies and aggregating model parameters received from clients. The training process follows the standard FL protocol (McMahan et al., 2017). For computational efficiency, we adopt Low-Rank Adaption (LoRA) (Hu et al., 2022) as the fine-tuning method, which involves tuning additional parameters $\Delta\phi$ while keeping the pre-trained LLM’s parameters $\phi$ frozen. In the initial training phase, the server dispatches $\phi$ and $\Delta\phi$ to each client for $\mathcal{R}$ training rounds. In the $t$-th round, the server sends the aggregated $\Delta\phi^{t}$ to clients. Clients use $\Delta\phi^{t}$ to update their local LoRA parameters $\Delta\phi_{k}^{t}$ and conduct instruction tuning based on augmented instruction datasets $D_{k}$. Subsequently, the clients return $\Delta\phi_{k}^{t+1}$ to the server. The server then aggregates $\{\Delta\phi_{k}^{t+1}\mid k=1,\dots,N\}$ to obtain $\Delta\phi^{t+1}$ for the next round.
@@ 3 Problem Formulation
For better understanding, we list the frequently used notation in Appendix A.11. The objective of FedDIT is to enhance the domain-specific performance of LLMs through FL without sharing private data (Ye et al., 2024b; Zhang et al., 2024b). Under the FL framework, suppose we have $N$ clients, where each client $c_{k}$ has a local dataset $D_{k}^{l}$ with its size $N_{k}^{l}$, and an augmented dataset $D_{k}^{g}$ from the server public dataset $D^{p}$, respectively. Due to constraints such as memory, computational overhead, and maximum tolerated training time, client $c_{k}$ can accept at most $N_{k}^{p}$ public instructions. Denote the augmentation strategy as $\Lambda$, through which the server performs instruction augmentation on the public dataset. If $\Lambda$ is null, it indicates that clients conduct FedDIT solely based on their local private data, which may lead to performance degradation. Conversely, $\Lambda$ may be classified into two categories: (1) focusing exclusively on the client’s own local data distribution or (2) considering the cross-client data distribution. In conclusion, the global objective of FedDIT is defined as follows:
@@ 4.1 Optimization Problem
Suppose there exists a metric space $\mathcal{X}$, a set of cross-client local instruction embeddings $\mathcal{E}\in\mathcal{X}$, and a set of cluster centers $\mathcal{P}\in\mathcal{X}$. Each client $c_{k}$ has maximum $\xi$ clusters obtained by k-means algorithm (Wu, 2012). In the federated setting, communication cost is always a critical factor. Consequently, we formulate the optimization problem to include communication costs as follows:
@@ 4.4 Discussions
Computation. Through greedy client center selection, FedDCA solves the optimization problem defined in Eq. 3 in the polynomial time and is quite efficient. On the client side, the k-means algorithm (Wu, 2012) is $\mathcal{O}(|D_{k}^{l}|)$. On the server side, as mentioned in Section 4.1, the time complexity of greedy client center selection is $\mathcal{O}\left(N\left(N\xi+\xi\log\xi\right)\right)$. Specifically, for each client, the similarity computation between selected client center set $\mathcal{P}$ and client $c_{k}$’s cluster $\mathcal{C}_{k}$ is $\mathcal{O}\left(N\xi\right)$, and the sorting process is $\mathcal{O}\left(\xi\log\xi\right)$. For domain data retrieval, given that the sorting algorithm is $\mathcal{O}\left(\mathcal{N}\log\mathcal{N}\right)$, where $\mathcal{N}$ is the size of the public dataset. Thus the time complexity is $\mathcal{O}\left(N(\mathcal{N}\log\mathcal{N})\right)$. In addition, the heterogeneous encoder setting FedDCA further reduces the computation overhead on the client side.
@@
Privacy. The proposed method, FedDCA, does not violate the client’s privacy. Compared with other FedIT methods (Zhang et al., 2024b; Ye et al., 2024b), the difference lies in the greedy client center selection. In this stage, the client only uploads the cluster center to the server, which is the average of embeddings to its cluster. In addition, the potential privacy leakage can be further avoided through homomorphic encryption (Acar et al., 2018), which allows the server to directly compute on ciphertext for matrix multiplication for dense retrieval.
@@ 5.1 Experimental Setup
Baselines. We compare FedDCA and FedDCA with the following baselines: 1) Unaugmented methods, including zero-shot inference and FedIT, which are composed of four widely used FL methods, including FedAvg (McMahan et al., 2017), FedProx (Li et al., 2020), SCAFFOLD (Karimireddy et al., 2020) and FedAvgM (Hsu et al., 2019). 2) Augmented methods, which include random sampling, direct retrieval, LESS (Xia et al., 2024), and Self-Instruct (Wang et al., 2022). Additionally, we report FedDCA’s performance with different FL strategies in Table 2. Specifically, zero-shot inference shows the performance of the pre-trained LLM without FedDIT, which gives the lower performance bound. Direct retrieval is described in Appendix A.3. Self-Instruct augments instruction in a generative way through GPT-3.5-turbo (Sun et al., 2023). Based on the prompt provided by Self-Instruct, we define the prompts for generating instructions and responses as shown in Appendix A.7.
@@ 5.2 Performance Analysis
Efficiency. We see that direct retrieval is the best baseline in the average performance. However, it does not consider cross-client domain coverage, resulting in an overlapping retrieved data distribution. Self-Instruct (Wang et al., 2022) represents the upper limit performance of the generation-based method, which is restricted to the high expense of cost, limited API call frequency, and heavy quality screening process. On the other hand, FewFedPIT (Zhang et al., 2024b) attempts to augment local data by leveraging the pre-trained Llama2 model to perform self-instructed generation during training. However, two concerns exist: 1) The pre-trained LLM cannot provide effective, stable, and high-quality instruction generation. 2) Generating instructions locally is very costly regarding both time and computational resources. Overall, generating instructions self-instructively is not a satisfactory method for the FedDIT scenario. LESS (Xia et al., 2024) represents the augmentation methods through gradient feature retrieval. However, its performance is underwhelming. Compounding this issue, LESS presupposes access to the validation set, which is not always available, and necessitates an initial warmup on the public dataset before calculating gradients for both the public and validation data. This process becomes computationally burdensome with large public datasets. In conclusion, FedDCA achieves better performance while requiring lower computational resources and a more relaxed training condition, which shows its efficiency.
@@ 5.4 Privacy Analysis
Next, we evaluate the privacy-preserving capability of different ratios of public data against memory extraction attacks (Carlini et al., 2021; Zhang et al., 2024a), which utilizes the autoregression nature of LLM to extract information from the memory (Xu et al., 2024).
@@
Retrieval Num. Code Med. Fin. Math.
@@ A.1 Related Work
Federated Instruction Tuning. Instruction tuning has been widely applied across various application areas of large language models (LLM), serving as a key technique to enhance the capabilities and controllability of LLM (Zhang et al., 2023d; Wei et al., 2022). Recently, federated instruction tuning (FedIT) has emerged as an effective strategy for the distributed optimization of LLMs, leveraging federated learning (FL) protocols to improve the handling of privacy-sensitive tasks in real-world scenarios. So far, several FedIT frameworks (Ye et al., 2024b; a) have been established to evaluate the effectiveness of FedIT across multiple datasets, tasks, and FL methods. While these platforms provide a foundation for research, they have not yet introduced more complex federated algorithms and deeply investigate the challenging problems and factors affecting FedIT, which are crucial for advancing this field.
@@
PrivateLoRA (Wang et al., 2023b) addresses privacy and efficiency issues by exploiting the low-rank properties of residual activations to reduce communication costs, significantly lowering the communication overhead through collaborative computation between the server and clients while effectively maintaining the privacy of local data. While FewFedPIT (Zhang et al., 2024b) focuses on the few-shot learning setting in FedIT, using self-generated data by pre-trained LLMs locally to mitigate the paucity of data and first discussing memory extraction attacks within FedIT.
@@
Domain Instruction Augmentation. In the real world, there is an urgent need for training LLMs with specific functionalities (e.g., reasoning capabilities) or domain-specific LLMs (e.g., code (Nijkamp et al., 2023; Luo et al., 2024), medical (Zhang et al., 2023e), financial (Yang et al., 2023b; a; Zhang et al., 2023a; Wu et al., 2023), mathematical (Yue et al., 2024; Luo et al., 2023)). Existing works tend to use open-source domain-specific instruction tuning datasets for training. However, the target domain may not always have corresponding ready-made domain-specific instruction datasets. Even if they exist, these datasets are often limited in scale.
@@
Several studies investigate domain-specific instruction augmentation, which can be categorized into three aspects: 1) Reusing human-curated public datasets (Wang et al., 2023a; Zhang et al., 2023f; Jiao et al., 2023; Lee et al., 2024; Xia et al., 2024). For instance, Parrot (Jiao et al., 2023) enhances translation capabilities of LLMs in chat by converting bilingual sentences into instruction-following datasets. Furthermore, works like INSTA (Lee et al., 2024) and LESS (Xia et al., 2024) attempt efficient domain-specific instruction augmentation via dense retrieval. INSTA (Lee et al., 2024) uses instructions without responses for effective retrieval. LESS (Xia et al., 2024) assumes access to a validation set and uses the warmup LLM’s gradients of the train and validation sets for retrieval. 2) Using seed tasks for self-instruct (Luo et al., 2024; Wan et al., 2023): Explore-Instruct (Wan et al., 2023), for example, employs activate exploration to tree-model domain tasks from both depth and breadth, increasing seed instructions’ coverage of the domain, and subsequently generating broader in-domain instructions through self-instruct. 3) Scaling instructions from the web: Recent works (Yue et al., 2024; Zhou et al., 2024) highlight the immense potential of mining naturally occurring instructions from the internet. Compared to generated data, web-mined instructions exhibit less bias and greater diversity. MAmmoTH2 (Yue et al., 2024) firstly retrieves domain-relevant texts and employs a LLM to extract Q-A pairs, further refining them into instruction-response pairs. Jiuzhang3.0 (Zhou et al., 2024) distills GPT’s instruction generation capabilities into a smaller model and then uses it to generate instructions from the internet. In conclusion, models fine-tuned with augmented instructions have shown promising domain-specific capabilities.